A company's profit is 15% lower in june than in may, due to unforeseen changes in weather. The total profit for both months together is $32,375
a. use an algebraic equations with two variables (x,y)
b. solve for the profits for each month separately: may and june

Answer by stanbon(75887)

A company's profit is 15% lower in june than in may, due to unforeseen changes in weather. The total profit for both months together is $32,375
a. use an algebraic equations with two variables (x,y)
Let x = june and y = may
---
x = 0.85y
x + y = 32375
--------------
b. solve for the profits for each month separately: may and june
Substitute for "x" and solve for "y"::
0.85y + y = 32375
1.85y = 32375
y = $17500
----
x = 0.85y = 0.85*17500 = $14875
------------
Cheers,
Stan H
--------------